OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. X/2 + 2/3 = 1/4

OpenStudy (texaschic101):

LCD of 2,3, and 4 is 12. x/2 + 2/3 = 1/4 (multiply everything by 12 6x + 8 = 3 6x = 3 - 8 6x = -5 x = -5/6 check.. (-5/6)/2 + 2/3 = 1/4 (-5/6* 1/2) + 2/3 = 1/4 -5/12 + 2/3 = 1/4 -5/12 + 8/12 = 1/4 3/12 = 1/4 1/4 = 1/4 (correct)